# MiniLLM: Knowledge Distillation of Large Language Models

### v5

@@ 3.2 Results
Second, the Rouge-L scores show that MiniLLM produces the most precise responses that have high overlaps with the ground-truth responses. In some cases, especially on Vicuna, S-NI, and UnNI, student models reach even higher Rouge-L scores than the teacher models, which matches the observation in [born_again]. We conjecture that the standard teacher-forcing fine-tuning on $\mathcal{D}$ brings training-inference discrepancy to the teacher model, also known as exposure bias [exposure_bias]. On the contrary, MiniLLM is optimized with policy optimization methods, which samples responses from student models during training and thus alleviates exposure bias [cold]. We include further analysis on exposure bias in Section 3.3.
@@ 3.3 Analysis
Although it is intuitive that we can distill better student models from larger teacher models, [teaching_assistant] has shown that increasing the teacher models’ sizes does not guarantee the improvement of student models, sometimes even harming the distillation performance. It is not clear how MiniLLM works when we scale up the teacher models’ sizes. Therefore, we compare MiniLLM and SeqKD using teacher models with different sizes and fix the size of the student model. We present the results based on the GPT-2 family in Figure 5 and that based on the OPT family in Appendix C.2. We can see that MiniLLM constantly outperforms SeqKD, and the student model performance is positively correlated with the teacher model sizes. This shows the potential of our method to compress models with massive parameters.
@@
Language generation models trained to minimize forward KLD suffer from exposure bias [exposure_bias] caused by the discrepancy between teacher-forcing training and free-run generation. When training MiniLLM, the student model sees samples generated by itself, alleviating the training-inference mismatch [cold]. In Figure 6, we use the ExAccErr metric [eb_measure] defined in Appendix B.5 to measure the excess accumulated error due to exposure bias. The experiment is based on GPT-2-125M, with GPT-2-1.5B as the teacher, using Dolly as the test set. For each prompt, we sample 10 responses to reduce the variance. We can see that the ExAccErrs of the baselines continuously grow during generation, while MiniLLM has a much lower ExAccErr, and the error stops accumulating in long-text generation (>> 150 tokens).
@@
[gpt4] has shown that models trained with policy optimization are likely to be poorly calibrated. We test the calibration of MiniLLM and the KD baselines on two widely-used text classification datasets: SST2 [sst-2] and BoolQ [boolq], based on LLaMA-7B. We design zero-shot classification instructions (see Appendix B.2) and take the probability of the label words to compute the ECE scores [ece]. From Table 6, we observe that KD and SeqKD models are worse calibrated than the teacher model, which potentially explains their low performance on canonical benchmarks [false_imitate]. We suspect that minimizing forward KLD causes the models to push high probabilities to void regions of the target distribution, which leads to significant distribution differences between the student and the teacher (see the example in Figure 2). In contrast, MiniLLM focuses on accurately learning the major parts of the target distribution and narrows the ECE scores gap between the student and the teacher.
@@
We study the models’ performance when the golden response lengths belong to different ranges. In Figure 7, we illustrate the Rouge-L scores of different KD models against the SFT models on three S-NI subsets split by the length of the ground truth responses. We can see that all methods achieve low scores on prompts that expect short responses ($\leq 5$ tokens), probably because most responses in our training set are long sentences, introducing a distribution shift between training and evaluation [ITGPT4]. Furthermore, the output spaces of these prompts are relatively small, allowing the student model to cover most modes of the teacher, and thus reverse KLD and forward KLD have similar performance. For prompts with longer responses ($\geq 6$ tokens), the teacher distribution contains more modes than the students due to the complex output spaces, which shows the advantage of MiniLLM against standard KD models. Similar results on UnNI are shown in Appendix C.3.
@@
[gan_falling_short] has found that the model optimized by minimizing reverse KLD is likely to lose modes, which affects the generation diversity. We follow [cold] to discuss generation diversity from three aspects: (i) generating multiple distinct responses given a prompt. (ii) generating linguistically complex responses. (iii) the ability to generate contents that have high coverage of the real data distribution. For (i), we argue that for many NLP applications, generating one correct response is sufficient, especially for those scenarios demanding high truthfulness and reliability [halu_survey]. For (ii) and (iii), we report the responses’ distinct 4-gram proportion and the language modeling loss on the test sets in Table 7, using the base models from the LLaMA family (see Appendix B.4 for more details) . We can see that MiniLLM preserves the distinct 4-gram proportion in the generated responses and language modeling loss on the test set.
@@ 3.4 Ablation Studies on Optimization Strategies
We evaluate the effectiveness of the three strategies proposed to stabilize and accelerate optimization in Section 2.2 by distilling a GPT-2-125M model from the GPT-2-1.5B model. More ablation studies can be found in Appendix C.4. In Table 8, we report the best Rouge-L scores on the validation set of each run and the evaluation results of the corresponding checkpoints. We also plot the reverse KLD between the student and the teacher during training in Figure 8, where the curves are smoothed by 32 steps. We can see that Teacher-Mixed Sampling and Length Normalization works for stabilizing training. Although the reverse KLDs also decrease without these strategies, we find that the models quickly learn to generate repeated, short, or meaningless strings that have high probabilities in the teacher distribution (see examples in Appendix D), which is known as reward hacking [reward_hacking]. This also leads to the low generation performance in Table 8. From Figure 8, we also observe that the Single-Step Decomposition effectively reduces the variance of the training process, which also results in higher scores on the validation and test sets.
@@ 4 Related Work
Large language models (LLMs; gpt3, lamda, palm, gpt4, palm2) have shown superior performance by solving various NLP tasks in a generative manner. Recent works apply instruction tuning [flan, t0, flan-t5] or learning from human feedback [instruct-gpt, hhh] to improve the alignment of LLMs with humans further and create general AI assistants [chatgpt, bard]. There are also efforts to build open-source LLMs [opt, llama, pythia] to facilitate research and industry development. Although appealing, the broad capacities of LLMs usually emerge with large model sizes [scaling_law, emergent] that require massive computational resources. Therefore, model compression is critical for the practical deployment and further research of LLMs.
@@
Knowledge distillation (KD; kd), as a widely used model compression technique, aims at training a student model with the guidance of a teacher model [policy_kd, distilbert, kdsurvey]. In the NLP community, many works apply KD to text classification tasks by mimicking the teacher model’s output distribution [lightpaff, mixkd, perturb_kd], hidden states [tinybert, bert-pkd], or attention scores [minilm, minilmv2]. For text generation, the standard KD method is to approximately minimize the forward KLD between the student’s and the teacher’s generation distribution by using the teacher’s output at each time step as supervision [distilbert] or direct training on the teacher’s generated texts [skd, alpaca, vicuna, ITGPT4]. In this paper, we minimize the reverse KLD, which is more suitable for LLMs when the teacher distribution is available. Concurrent works [gkd, f-div-kd] also explore more the distribution discrepancy metrics in KD.
@@
The distribution discrepancy metrics play a significant role in training text generation models. The forward KLD is widely used due to its simplicity when derived as the Maximum Likelihood Estimate (MLE) objective [zhang2018minimum]. However, previous works show that minimizing forward KLD leads to zero-forcing behavior where models try to cover all modes of the target distribution and sacrifice the accuracy of major modes [huszar2015not]. Some works resort to using other metrics to remedy this problem, such as reverse KLD [smart], Total Variation Distance [tailor], and Optimal Transport [ot_for_gen]. Our paper tackles this problem under the scenario of knowledge distillation for LLMs.
@@ A.1 A Perspective of MiniLLM from Inverse Reinforcement Learning
In Section 2.1, we formulate KD as an optimization problem of minimizing the discrepancy between the teacher distribution and the student distribution and finally reach the objective of minimizing reverse KLD. Alternatively, we can also regard KD as training the student model with the teacher model’s guidance, which resembles an agent learning from the feedback from an environment. Following [cold], we treat token generation as a Markov Decision Process. At each time step $t$, the student model chooses an action (token) $y_{t}$ from the action space (vocabulary) $V$ conditioning on the state (prefix) $(\bm{y}_{<t},\bm{x})$ based on the policy (generation probability) $q_{\theta}(y_{t}|\bm{y}_{<t},\bm{x})$.
From this perspective, standard KD corresponds to behavior cloning (BC; behavior_cloning) in imitation learning [imitation_learning]. However, BC is known to under-perform Inverse Reinforcement Learning (IRL; maxent_irl), another imitation learning method that first recovers a reward model from the environment demonstrations and then trains the policy with the reward using policy optimization algorithms [policy_gradient, ppo]. Therefore, in the KD scenario, we seek to first induce a reward $r(y_{t},(\bm{y}_{<t},\bm{x}))$ from the environment (the teacher model) and then train the student model to maximize the reward as the objective. We take the maximum-entropy inverse reinforcement learning framework [maxent_irl, scalable_irl] and thus the Q-function $Q(y_{t},(\bm{y}_{<t},\bm{x}))$ in the environment satisfies the soft Bellman Equation:
@@
We follow [irl_gen] to parameterize the Q-function as $Q(y_{t},(\bm{y}_{<t},\bm{x}))=f(y_{t},(\bm{y}_{<t},\bm{x}))$ and assume $\gamma=1$, where $f(y_{t},(\bm{y}_{<t},\bm{x}))$ is the output logits of the teacher model [^4] The teacher model’s distribution satisifies $p(y_{t}|\bm{y}_{<t},\bm{x})=\frac{\operatorname{exp}(f(y_{t},(\bm{y}_{<t},\bm{x})))}{\sum_{y^{\prime}\in V}\operatorname{exp}(f(y^{\prime},(\bm{y}_{<t},\bm{x})))}$.. Then, the reward is given by:
@@
To maximize the reward, we apply maximum-entropy reinforcement learning [max_ent_rl], whose learning objective is
@@ A.2 Derivation of Equation 2
We compute the gradient of $\mathcal{L}(\theta)=\operatorname{KL}[q_{\theta}||p]$ with respect to $\theta$ using the Policy Gradient Theorem [policy_gradient]:
@@ B.1 Training Details
Our baselines include SFT w/o KD, KD, and SeqKD. For models with less than 1.3B parameters, we search for the learning rates in [5e-4, 1e-4, 5e-5], the batch sizes in [32, 64], and train these models for 20 epochs. For other models, we search for the learning rate in [5e-5, 1e-5, 5e-6], the batch sizes in [32, 64], and train these models for 10 epochs. For KD, we follow [lightpaff] to mix the distillation loss with the language modeling loss on the ground truth responses by a mixture rate of 0.5. The checkpoints of each baseline are selected by the Rouge-L [rouge] scores on the validation set because, as stated in previous works [super-natural-instructions, instruct-gpt], we also find that Rouge-L is better correlated with human judgments.
@@
As stated in Section 2.3, training of MiniLLM has two phases which is similar to Reinforcement Learning from Human Feedback (RLHF;instruct-gpt).
@@
  • Phase 2: We continuously train the model from Phase 1 as described in Algorithm 1 using a learning rate 5e-6, a mini-batch size $64$ in all cases. The training and validation set are same as in Phase 1. Similar to [instruct-gpt], we collect 256 sentences at once and adopt 4 inner epochs when doing the policy optimization. The clipping rate $\epsilon$ is set to 0.2, and the max length of the model is 512. We use temperature = 1 when sampling from $q_{\theta}$. We train the model for 5000 steps and select the final checkpoint using the Rouge-L score on the validation set. Our experiments are based on the NVIDIA V100 32G GPUs. Distilling LLaMA-7B from LLaMA-13B takes less than 10 ours on 16 GPUs.
@@ B.2 Automatic Evaluation Details
During the evaluation, we sample the responses from each model using temperature = 1, a max-length limit of 512, and random seeds [10, 20, 30, 40, 50]. Similar to [alpaca], we adopt a prompt wrapper shown in Figure 9 to convert each instruction-response pair to a sentence. For the GPT-4 feedback, we apply the prompt in Figure 10 and set the temperature = 0.7. For the classification tasks in the “Calibration” paragraph of Section 3.3, we prompt the model to do zero-shot text classification with the templates in Figure 11 and 12.
@@ B.3 Human Evaluation Details
Following [ITGPT4], we use SelfInst [self_inst] to perform human evaluation. We randomly sampled 50 prompts because we found that more prompts do not affect the results much. We ask the annotators to compare the responses generated by the baseline models with MiniLLM and decide which response is preferred or neither of them is significantly better. Note that which model the responses come from is invisible to the annotators. The interface presented to annotators is shown in Figure 13.
@@ B.4 Details About Generation Diversity Metrics
  • “Dist-4” is a fraction: $N/C$ , where $N$ is the number of the distinct 4-grams in the generated responses and $C$ is the total number of 4-grams. It is a widely used metric to measure the generation diversity of a language model [dist4]. The $(N/C)$s on the Dolly test set across 5 random seeds are shown in Table 5. Table 7 reports the average values across the 5 random seeds.
@@ B.5 Exposure Bias Analysis
Following [eb_measure], we compute the ExAccErr with the following formula:

### v6

@@ 3.2 Results
Second, the Rouge-L scores show that MiniLLM produces the most precise responses that have high overlaps with the ground-truth responses. In some cases, especially on Vicuna, S-NI, and UnNI, student models reach even higher Rouge-L scores than the teacher models, which matches the observation in [19]. We conjecture that the standard teacher-forcing fine-tuning on $\mathcal{D}$ brings training-inference discrepancy to the teacher model, also known as exposure bias [5]. On the contrary, MiniLLM is optimized with policy optimization methods, which samples responses from student models during training and thus alleviates exposure bias [52]. We include further analysis on exposure bias in Section 3.3.
@@ 3.3 Analysis
Although it is intuitive that we can distill better student models from larger teacher models, [46] has shown that increasing the teacher models’ sizes does not guarantee the improvement of student models, sometimes even harming the distillation performance. It is not clear how MiniLLM works when we scale up the teacher models’ sizes. Therefore, we compare MiniLLM and SeqKD using teacher models with different sizes and fix the size of the student model. We present the results based on the GPT-2 family in Figure 5 and that based on the OPT family in Appendix C.2. We can see that MiniLLM constantly outperforms SeqKD, and the student model performance is positively correlated with the teacher model sizes. This shows the potential of our method to compress models with massive parameters.
@@
Language generation models trained to minimize forward KLD suffer from exposure bias [5] caused by the discrepancy between teacher-forcing training and free-run generation. When training MiniLLM, the student model sees samples generated by itself, alleviating the training-inference mismatch [52]. In Figure 6, we use the ExAccErr metric [3] defined in Appendix B.5 to measure the excess accumulated error due to exposure bias. The experiment is based on GPT-2-125M, with GPT-2-1.5B as the teacher, using Dolly as the test set. For each prompt, we sample 10 responses to reduce the variance. We can see that the ExAccErrs of the baselines continuously grow during generation, while MiniLLM has a much lower ExAccErr, and the error stops accumulating in long-text generation (>> 150 tokens).
@@
[50] has shown that models trained with policy optimization are likely to be poorly calibrated. We test the calibration of MiniLLM and the KD baselines on two widely-used text classification datasets: SST2 [62] and BoolQ [17], based on LLaMA-7B. We design zero-shot classification instructions (see Appendix B.2) and take the probability of the label words to compute the ECE scores [47]. From Table 6, we observe that KD and SeqKD models are worse calibrated than the teacher model, which potentially explains their low performance on canonical benchmarks [22]. We suspect that minimizing forward KLD causes the models to push high probabilities to void regions of the target distribution, which leads to significant distribution differences between the student and the teacher (see the example in Figure 2). In contrast, MiniLLM focuses on accurately learning the major parts of the target distribution and narrows the ECE scores gap between the student and the teacher.
@@
We study the models’ performance when the golden response lengths belong to different ranges. In Figure 7, we illustrate the Rouge-L scores of different KD models against the SFT models on three S-NI subsets split by the length of the ground truth responses. We can see that all methods achieve low scores on prompts that expect short responses ($\leq 5$ tokens), probably because most responses in our training set are long sentences, introducing a distribution shift between training and evaluation [53]. Furthermore, the output spaces of these prompts are relatively small, allowing the student model to cover most modes of the teacher, and thus reverse KLD and forward KLD have similar performance. For prompts with longer responses ($\geq 6$ tokens), the teacher distribution contains more modes than the students due to the complex output spaces, which shows the advantage of MiniLLM against standard KD models. Similar results on UnNI are shown in Appendix C.3.
@@
[10] has found that the model optimized by minimizing reverse KLD is likely to lose modes, which affects the generation diversity. We follow [52] to discuss generation diversity from three aspects: (i) generating multiple distinct responses given a prompt. (ii) generating linguistically complex responses. (iii) the ability to generate contents that have high coverage of the real data distribution. For (i), we argue that for many NLP applications, generating one correct response is sufficient, especially for those scenarios demanding high truthfulness and reliability [31]. For (ii) and (iii), we report the responses’ distinct 4-gram proportion and the language modeling loss on the test sets in Table 7, using the base models from the LLaMA family (see Appendix B.4 for more details) . We can see that MiniLLM preserves the distinct 4-gram proportion in the generated responses and language modeling loss on the test set.
@@ 3.4 Ablation Studies on Optimization Strategies
We evaluate the effectiveness of the three strategies proposed to stabilize and accelerate optimization in Section 2.2 by distilling a GPT-2-125M model from the GPT-2-1.5B model. More ablation studies can be found in Appendix C.4. In Table 8, we report the best Rouge-L scores on the validation set of each run and the evaluation results of the corresponding checkpoints. We also plot the reverse KLD between the student and the teacher during training in Figure 8, where the curves are smoothed by 32 steps. We can see that Teacher-Mixed Sampling and Length Normalization works for stabilizing training. Although the reverse KLDs also decrease without these strategies, we find that the models quickly learn to generate repeated, short, or meaningless strings that have high probabilities in the teacher distribution (see examples in Appendix D), which is known as reward hacking [61]. This also leads to the low generation performance in Table 8. From Figure 8, we also observe that the Single-Step Decomposition effectively reduces the variance of the training process, which also results in higher scores on the validation and test sets.
@@ 4 Related Work
Large language models (LLMs; 9, 67, 14, 50, 2) have shown superior performance by solving various NLP tasks in a generative manner. Recent works apply instruction tuning [75, 58, 15] or learning from human feedback [51, 4] to improve the alignment of LLMs with humans further and create general AI assistants [49, 21]. There are also efforts to build open-source LLMs [82, 69, 6] to facilitate research and industry development. Although appealing, the broad capacities of LLMs usually emerge with large model sizes [35, 76] that require massive computational resources. Therefore, model compression is critical for the practical deployment and further research of LLMs.
@@
Knowledge distillation (KD; 26), as a widely used model compression technique, aims at training a student model with the guidance of a teacher model [56, 57, 33]. In the NLP community, many works apply KD to text classification tasks by mimicking the teacher model’s output distribution [63, 41, 81], hidden states [34, 64], or attention scores [72, 71]. For text generation, the standard KD method is to approximately minimize the forward KLD between the student’s and the teacher’s generation distribution by using the teacher’s output at each time step as supervision [57] or direct training on the teacher’s generated texts [36, 66, 13, 53]. In this paper, we minimize the reverse KLD, which is more suitable for LLMs when the teacher distribution is available. Concurrent works [1, 77] also explore more the distribution discrepancy metrics in KD.
@@
The distribution discrepancy metrics play a significant role in training text generation models. The forward KLD is widely used due to its simplicity when derived as the Maximum Likelihood Estimate (MLE) objective [80]. However, previous works show that minimizing forward KLD leads to zero-forcing behavior where models try to cover all modes of the target distribution and sacrifice the accuracy of major modes [29]. Some works resort to using other metrics to remedy this problem, such as reverse KLD [32], Total Variation Distance [30], and Optimal Transport [39]. Our paper tackles this problem under the scenario of knowledge distillation for LLMs.
@@ A.1 A Perspective of MiniLLM from Inverse Reinforcement Learning
In Section 2.1, we formulate KD as an optimization problem of minimizing the discrepancy between the teacher distribution and the student distribution and finally reach the objective of minimizing reverse KLD. Alternatively, we can also regard KD as training the student model with the teacher model’s guidance, which resembles an agent learning from the feedback from an environment. Following [52], we treat token generation as a Markov Decision Process. At each time step $t$, the student model chooses an action (token) $y_{t}$ from the action space (vocabulary) $V$ conditioning on the state (prefix) $(\bm{y}_{<t},\bm{x})$ based on the policy (generation probability) $q_{\theta}(y_{t}|\bm{y}_{<t},\bm{x})$.
From this perspective, standard KD corresponds to behavior cloning (BC; 68) in imitation learning [16]. However, BC is known to under-perform Inverse Reinforcement Learning (IRL; 85), another imitation learning method that first recovers a reward model from the environment demonstrations and then trains the policy with the reward using policy optimization algorithms [65, 59]. Therefore, in the KD scenario, we seek to first induce a reward $r(y_{t},(\bm{y}_{<t},\bm{x}))$ from the environment (the teacher model) and then train the student model to maximize the reward as the objective. We take the maximum-entropy inverse reinforcement learning framework [85, 11] and thus the Q-function $Q(y_{t},(\bm{y}_{<t},\bm{x}))$ in the environment satisfies the soft Bellman Equation:
@@
We follow [25] to parameterize the Q-function as $Q(y_{t},(\bm{y}_{<t},\bm{x}))=f(y_{t},(\bm{y}_{<t},\bm{x}))$ and assume $\gamma=1$, where $f(y_{t},(\bm{y}_{<t},\bm{x}))$ is the output logits of the teacher model [^4] The teacher model’s distribution satisifies $p(y_{t}|\bm{y}_{<t},\bm{x})=\frac{\operatorname{exp}(f(y_{t},(\bm{y}_{<t},\bm{x})))}{\sum_{y^{\prime}\in V}\operatorname{exp}(f(y^{\prime},(\bm{y}_{<t},\bm{x})))}$.. Then, the reward is given by:
@@
To maximize the reward, we apply maximum-entropy reinforcement learning [23], whose learning objective is
@@ A.2 Derivation of Equation 2
We compute the gradient of $\mathcal{L}(\theta)=\operatorname{KL}[q_{\theta}||p]$ with respect to $\theta$ using the Policy Gradient Theorem [65]:
@@ B.1 Training Details
Our baselines include SFT w/o KD, KD, and SeqKD. For models with less than 1.3B parameters, we search for the learning rates in [5e-4, 1e-4, 5e-5], the batch sizes in [32, 64], and train these models for 20 epochs. For other models, we search for the learning rate in [5e-5, 1e-5, 5e-6], the batch sizes in [32, 64], and train these models for 10 epochs. For KD, we follow [63] to mix the distillation loss with the language modeling loss on the ground truth responses by a mixture rate of 0.5. The checkpoints of each baseline are selected by the Rouge-L [42] scores on the validation set because, as stated in previous works [74, 51], we also find that Rouge-L is better correlated with human judgments.
@@
As stated in Section 2.3, training of MiniLLM has two phases which is similar to Reinforcement Learning from Human Feedback (RLHF;51).
@@
  • Phase 2: We continuously train the model from Phase 1 as described in Algorithm 1 using a learning rate 5e-6, a mini-batch size $64$ in all cases. The training and validation set are same as in Phase 1. Similar to [51], we collect 256 sentences at once and adopt 4 inner epochs when doing the policy optimization. The clipping rate $\epsilon$ is set to 0.2, and the max length of the model is 512. We use temperature = 1 when sampling from $q_{\theta}$. We train the model for 5000 steps and select the final checkpoint using the Rouge-L score on the validation set. Our experiments are based on the NVIDIA V100 32G GPUs. Distilling LLaMA-7B from LLaMA-13B takes less than 10 ours on 16 GPUs.
@@ B.2 Automatic Evaluation Details
During the evaluation, we sample the responses from each model using temperature = 1, a max-length limit of 512, and random seeds [10, 20, 30, 40, 50]. Similar to [66], we adopt a prompt wrapper shown in Figure 9 to convert each instruction-response pair to a sentence. For the GPT-4 feedback, we apply the prompt in Figure 10 and set the temperature = 0.7. For the classification tasks in the “Calibration” paragraph of Section 3.3, we prompt the model to do zero-shot text classification with the templates in Figure 11 and 12.
@@ B.3 Human Evaluation Details
Following [53], we use SelfInst [73] to perform human evaluation. We randomly sampled 50 prompts because we found that more prompts do not affect the results much. We ask the annotators to compare the responses generated by the baseline models with MiniLLM and decide which response is preferred or neither of them is significantly better. Note that which model the responses come from is invisible to the annotators. The interface presented to annotators is shown in Figure 13.
@@ B.4 Details About Generation Diversity Metrics
  • “Dist-4” is a fraction: $N/C$ , where $N$ is the number of the distinct 4-grams in the generated responses and $C$ is the total number of 4-grams. It is a widely used metric to measure the generation diversity of a language model [40]. The $(N/C)$s on the Dolly test set across 5 random seeds are shown in Table 5. Table 7 reports the average values across the 5 random seeds.
@@ B.5 Exposure Bias Analysis
Following [3], we compute the ExAccErr with the following formula: